OpenStudy (anonymous):

Write \(log_8~3\) as a logarithm of base 5...?

OpenStudy (anonymous):

@mathstudent55 explain? :)

OpenStudy (mrdoe):

use the change of base formula, its been a long time since ive had to use it so i don't know how much help i can be past that, but you can google it.

OpenStudy (mrdoe):

heres a good site: http://www.sosmath.com/algebra/logs/log4/log43/log43.html

OpenStudy (anonymous):

i need someone to explain it..i dont get how to do it and google isnt helping its only causing more confusion..

OpenStudy (anonymous):

that didnt help :\

OpenStudy (anonymous):

@ChmE

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

Similar Example: \[\Large \log_{7}(14) = \frac{\log_{5}(14)}{\log_{5}(7)}\] Since in general \[\Large \log_{b}(x) = \frac{\log_{y}(x)}{\log_{y}(b)}\] for any positive real number y.

OpenStudy (mrdoe):

so it would be log8(3)=\[\frac{ \log5(3) }{ \log5(8)}\] can you take it from there?

OpenStudy (anonymous):

OHHH....thank you soo much